There are 19 animals in the field. Some are goats and some are ducks. There are 72 legs in all. How many of each animal are in the field?

Answer by ikleyn(52787)

.
There are 19 animals in the field. Some are goats and some are ducks.
There are 72 legs in all. How many of each animal are in the field?
~~~~~~~~~~~~~~

Let x be the number of goats;  then the number of ducks is (19-x).


There is no need to remind that the ducs have two legs, while the goats have four legs.


So we write an equation for the total of legs


    4x + 2*(19-x) = 72.


Simplify and find x


    4x + 38 - 2x = 72

    4x - 2x = 72 - 38

      2x    =   34

       x    =   34/2 = 17.


ANSWER.  There are 17 goats and 19-17 = 2 ducks.


CHECK.   17*4 + 2*2 = 68 + 4 = 72  total legs.   ! Correct !

Solved.

--------------

To see many other similar  (and different)  solved problems,  look into the lessons
    - Problem on two-wheel and three-wheel bicycles
    - Problem on animals at a farm
    - Problem on pills in containers
    - What type of problems are these?
at this site.